Flying over the western states with mountainous terrain in a small aircraft is 40% riskier than flying over similar distances in flatter portions of the nation, according to a General Accounting Office study completed in response to a congressional request. The accident rate for small aircraft in the 11 mountainous western states is 1.5 accidents per 100,000 flight operations.

(a) Let r = number of accidents for a given number of operations. Explain why the Poisson distribution would be a good choice for the probability distribution of r.

Frequency of accidents is a rare occurrence. It is reasonable to assume the events are independent.

Frequency of accidents is a common occurrence. It is reasonable to assume the events are dependent.  

  Frequency of accidents is a rare occurrence. It is reasonable to assume the events are dependent.

Frequency of accidents is a common occurrence. It is reasonable to assume the events are independent.



(b) Find the probability of no accidents in 100,000 flight operations. (Use 4 decimal places.)


(c) Find the probability of at least 6 accidents in 200,000 flight operations. (Use 4 decimal places.)
Answer #1

Frequency of accidents is a rare occurrence. It is reasonable to assume the events are independent.

b)

probability = P(X=0)= {ex/x!}= 0.2231

c)

probability = P(X>=6)= 1-P(X<=5)= 1-∑x=0x-1   {ex/x!}= 0.0045